Solve the following LPP using dual simplex method: Max Z= -3x1- x2,

Subject to: x1+x2 ≥ 1, 2x1+3x2 ≥ 2, x1,x2 ≥ 0.


1
Expert's answer
2021-07-01T13:59:24-0400

Solution.

"-3x_1-x_2\to max,"

"x_1+x_2\geq 1,\newline
2x_1+3x_2\geq 2."

We replace signs in restrictions and we enter additional variables x3, x4.

Initial simplex table, where x3 and x4 are basic variables.



So, as "|b|_{max}=|-2|" is in the 2nd line, and the maximum modulo element in the 2nd line is -3, x2 is a new basic variable.

Updated table

"|b|_{max}=|-\frac{1}{3}|"

is in 1 row, and the largest element modulo "-\frac{1}{3}" is in 1 column, then x1 is a new basic variable.

Updated table.



Let's calculate and create a simplex table with deltas.

"\u2206_i=C_1\u2022a_{1i}+C_2\u2022a_{2i}-C_i"


"\u2206_4=-2<0," therefore, the plan is suboptimal.

Take the new base variable x4 instead of x1.

We will have the following table, as well as a table with deltas.


"\u2206_i>0," therefore, the plan (0,1,0,1) is optimal.

So,

"F_{max}=-3\u20220+(-1)\u20221+0\u20220+0\u20221=-1."

Answer. -1.